OpenStudy (anonymous):

William bought fourteen bags of candy at the store for a total price of $46.24. He bought some bags of candy that cost $2.93 each and the rest cost $3.51 each. How many bags of the $2.93 candy did William buy?

OpenStudy (theloshua):

no, thats not the right set up for that question, since it specifically said He bought 14 bags of candy

OpenStudy (anonymous):

what should i do then? @Theloshua

OpenStudy (a1234):

I knew something was wrong with my equation... This one better be right :) x + y = 14 2.93x + 3.51y = 46.24 Can you solve it?

OpenStudy (a1234):

Do you know how to solve systems of equations?

OpenStudy (anonymous):

my answer doesn't look quite right?

OpenStudy (a1234):

What is it :) ?

OpenStudy (anonymous):

x 1.197952 y 15.784983 lol

OpenStudy (anonymous):

Solve the following for c. \[2.93 c + 3.51 (14 - c) = 46.24\]

OpenStudy (a1234):

x + y = 14 2.93x + 3.51y = 46.24 x represents 2.93 Solve for x x = 14 - y 2.93(14 - y) + 3.51y = 46.24

OpenStudy (anonymous):

c = 5

OpenStudy (anonymous):

5 bags of candy?

OpenStudy (anonymous):

Yes, 5 bags of the $2.93 each candy.

OpenStudy (anonymous):

thank you! @robtobey

OpenStudy (anonymous):

Your welcome.